For a particular quadratic function f(x)=ax^2+bx+c
f(-3)=-13 and f(2)=7. If a=-3, find b and c.
I know the points are (-3,-13) and (2,7).

Answer by Alan3354(69443)

For a particular quadratic function f(x)=ax^2+bx+c
f(-3)=-13 and f(2)=7. If a=-3, find b and c.
I know the points are (-3,-13) and (2,7).
--------------
f(x) = -3x^2 + bx + c
f(-3) = -3(-3)^2 + b(-3) + c = -13
-3b + c = 14 Eqn 1
-------
f(2) = -3*2^2 + b*2 + c = 7
2b + c = 19 Eqn 2
-3b + c = 14 Eqn 1
--------------------------- Subtract
5b = 5
b = 1 ****************
---
2b + c = 19 Eqn 2
c = 17 ***********